OpenStudy (mtepecz):

Write the expression in terms of sine only: -3cos2t+5sin2t

OpenStudy (owlcoffee):

like this? \(-3 \cos2t+5\sin2t\) ? or \(-3\cos^2t+5\sin^2t\) ?

OpenStudy (mtepecz):

The first one

imqwerty (imqwerty):

\[\cos2t=1-2\sin^{2}t\] just substitute

OpenStudy (jhannybean):

Remember that \(\cos(2t) = 1-2\sin^2(t)\)

OpenStudy (mtepecz):

@imqwerty @Jhannybean where do we get the cos (2t) from again?

imqwerty (imqwerty):

do u know this identity-\[\cos(A+B)=cosAcosB-sinAsinB\] ??

OpenStudy (mtepecz):

oh yeahhh

imqwerty (imqwerty):

yea try to replace A and B with t what will u get?

OpenStudy (mtepecz):

2t?

imqwerty (imqwerty):

yes and on the right hand side?

OpenStudy (baru):

i think the question is asking to write in the form Asin\((2 t-\phi)\)

OpenStudy (mtepecz):

yeah i have to write it with the sine formula

imqwerty (imqwerty):

yes we will do that but step wise :)

OpenStudy (jhannybean):

Seems to me everything just has to be in terms of sine at the end. Why not replace \(\cos(2t)\) in your problem with \(1-\sin^2(t)\) andf see where that takes you?

imqwerty (imqwerty):

@mtepecz u stuck somewhere?

OpenStudy (mtepecz):

I'm just confused on why its \[\cos( 2t)\] ansd not \[\cos ^{2}(2t)\]

imqwerty (imqwerty):

u mean like when we put t in this-> cos(A+B) then we get cos(2t) but y not cos^2(2t) right?

OpenStudy (mtepecz):

yeah

OpenStudy (mtepecz):

Yes

imqwerty (imqwerty):

ok now lets convert that \[\cos^2(t) \] into sine terms can u write it in terms of sine??

OpenStudy (mtepecz):

\[\cos ^{2}t=1-\sin ^{2}t\]

imqwerty (imqwerty):

yes correct :) now we put this value of cos^2(t) in our equation we get this-\[\cos(2t)=\cos^2(t)-\sin^2(t)\]now we put the value\[\cos(2t)=1-\sin^2(t)-\sin^2(t)\]simplifying\[\cos(2t)=1-2\sin^2(t)\] ok? :)

OpenStudy (mtepecz):

is that all or do we have to square root it?

imqwerty (imqwerty):

no no now we have an identity cos(2t)=1-2sin^2 (t) now u look ta ur question an see where u can use this identity

OpenStudy (mtepecz):

I honestly don't know

imqwerty (imqwerty):

its ok :) do u get that cos(2t)=1-2sin^2(t) ?

OpenStudy (mtepecz):

Yeah I get that, but what do we do next? Plug in numbers?

imqwerty (imqwerty):

ok now lets look at our original equation

imqwerty (imqwerty):

we have this\[-3\cos(2t)+5\sin(2t)\] we need to bring all the terms in sine can u tell me which term is not in sine

OpenStudy (mtepecz):

Would it be the right side?

imqwerty (imqwerty):

what do u mean by right side?

OpenStudy (mtepecz):

5sin 2t

imqwerty (imqwerty):

no :) well we need to bring all terms in sine form but 5sin(2t) is already in sine form but -3cos(2t) is in cosine form and we have to deal with it ok?

OpenStudy (mtepecz):

okay

imqwerty (imqwerty):

so we look at -3cos(2t) we see that this thing has got cos(2t) in it! and now we use our identity can u apply our identity in this?

ganeshie8 (ganeshie8):

Hey, as baru said, we could also write it as \[-3\cos 2t + 5\sin 2t = \sqrt{3^2+5^2} \sin(2t-\phi)\]

OpenStudy (mtepecz):

Is there going to be a square root at some point? I feel like what I'm doing is wrong

OpenStudy (mtepecz):

@imqwerty

imqwerty (imqwerty):

no ur not wrong are u feeling like -on one side its a linear term cos(2t) and on the other side we have sin^2 (t) which is a squared term so we shuld square root ?

OpenStudy (mtepecz):

yes

imqwerty (imqwerty):

haha don't worry then u can try an example if u want for example u take t=45degrees :) now u put it in our identity to check if it holds true

OpenStudy (mtepecz):

no

imqwerty (imqwerty):

it didn't hold true?

OpenStudy (mtepecz):

no I probably did it wrong

imqwerty (imqwerty):

check ur calculations once again

OpenStudy (mtepecz):

I'm sorry to have wasted your time, I'm tired and don't get it but thank you :)